What is the slope of the line y= -2x+5? I am not certain if what I graphed is right? please help.
Find the value of x at which the two lines y= 2x+1 and y+ -1/4x+3 intersect. I am confused. Thank you.

Answer by jim_thompson5910(35256)

the slope of the line y= -2x+5 is -2 (since y = mx+b and m = -2)

For the second one, start with y = 2x+1 and plug in y = -1/4x+3 to get

-1/4x+3 = 2x+1

and then solve for x.